A city has two subrubs subrub x anf subrub y over the past several years, the city has experienced a populatin shift from the city to the subrubs

1
Expert's answer
2022-01-17T07:12:56-0500

City (C)

Suburb x (X)

Suburb y (Y)



City (C)

.85

.07

.08



Suburb x (X)

.01

.96

.03



Suburb y (Y)

.01

.02

.97


In 20xo, the city had a population of 120,000, suburb x had a population of 80,000, and suburb by had a population of 50,000. Assuming that the population in the metropolitan area remains constant at 250,000 people.


Solution:

if data in table are shift of population in proportion of total population (250000 people), then:

for 20X2 year:

city; "120000+250000(-0.15+0.01)=85000" people

suburb x: "80000+250000(0.07-0.04)=87500" people

suburb y: "50000+250000(0.08+0.03)=77500" people